Support Services Meeting Notes
February 2, 2009
February 2, 2009
Present: Bill Covey, Will Chaney, Lori Driscoll, Rob Roberts, Laurie TaylorFacilities
- Hangar door downstairs - problem is the infrared sensors. They are out of alignment. We need to get a solid mount for the sensors.
- Electrical contractors are here. EHS inspected and there are standards that have not been met. The problems are being corrected.
- Rob is unable to access his email as a result of the UFAD migration. He has reported the issue via Grover.
- The white minivan had a tune-up recently. However, on a trip to South Florida last week, it started misbehaving. Should we take it to the dealer or to Bush Automotive? Rob is awaiting an answer from Fiscal Services. It is scheduled to go to St. Pete tomorrow and elsewhere on Thursday and Friday. It is not fit to go. A car may be rented if need be.
- Furniture was ordered in time for the projects in AFA, MSL and Education.
- An I-beam was run through the glass escalator panel (ascending escalator). The glass broke. There is a piece of plywood in there now.
- Problems with Illiad - Stephen Williams has been working to get it up and running in AD.
- Ariel is also not working. Lori reports that the tickets are being closed, but they are not checking to make sure all of the problems are resolved.
- Email is also not working for several people.
- Melissa Ware's vacated position - Angela is temporarily filling in. We can fill the post internally.
- University Library Committee to meet on Friday. No word on fines and fees.
- HSC/AS migration meeting via conference call - focused on Circulation policy, ERES and ILL. We will start up smaller focus groups and come up with a plan by February 17.
- AD migration - it went pretty well, considering how much work had to done.
- #UFLIB - Library default for the address book. We are currently working on service accounts - calendars and laptops.
- CNS wants to do work on the Education subnet.
- Cross-core routing application may come up with HSC.
- Greenstone - new drives to cost $2843. Will be used to replace three malfunctioning drives. The rest are for more drive power.
- Public folders have not been touched yet.
- The DLC is almost back up and running, post-AD migration. The Prime OCR machines are not working. Admin rights issue.
- Received drives for World News archive project.
- Drives for FCLA - met with them. We'll give them a couple of drives and see what they do.
- The way map ingest is done for PALMM is changing. The machine is being reprocessed.
- WinDev - want finding aides.
- Pete Hildebrand stopped by with stuff.
- President Machen wants the data for our 10% budget cutby March. The Dean wants it earlier so we can request funds for the HSC merger. She wants the data by this Friday. No more emails on the topic. OPS is to be zero-based. You should discuss it with staff in terms of positions, not names. There will be a group presentation at a Town Hall meeting on the budget. Let Bill know when your numbers are ready.
- Xerox replacement contract - 3rd party vendor - no solution yet.